OpenStudy (anonymous):

4. Solve the proportion using number sense. 2/8 X/48 A) 6 B) 12 C) 16 D) 8 5. Solve the proportion using cross products. 18/20 K/110 A) 99 B) 122 C) 3.3 D) 2.9 Yeah im really asking this xD im tired ok...

jimthompson5910 (jim_thompson5910):

2/8 = x/48 2*48 = 8*x ... cross multiply 96 = 8x 8x = 96 x = ???

OpenStudy (anonymous):

I know how to cross multiply...........

jimthompson5910 (jim_thompson5910):

ok, so what is x?

OpenStudy (anonymous):

its im going to the e.r. and i wanted my work for school to be done e.e